\begin{thebibliography}{8}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Agarwal et~al.(2014)Agarwal, Hsu, Kale, Langford, Li, and
  Schapire]{Agarwal2014}
Alekh Agarwal, Daniel Hsu, Satyen Kale, John Langford, Lihong Li, and Robert~E.
  Schapire.
\newblock Taming the monster: {A} fast and simple algorithm for contextual
  bandits.
\newblock In \emph{International Conference on Machine Learning (ICML)}, 2014.

\bibitem[Auer et~al.(1995)Auer, Cesa-Bianchi, Freund, and Schapire]{Auer1995}
Peter Auer, Nicolo Cesa-Bianchi, Yoav Freund, and Robert~E. Schapire.
\newblock Gambling in a rigged casino: The adversarial multi-armed bandit
  pproblem.
\newblock In \emph{Foundations of Computer Science (FOCS)}, 1995.

\bibitem[Cesa-Bianchi et~al.(1997)Cesa-Bianchi, Freund, Haussler, Helmbold,
  Schapire, and Warmuth]{cesa1997use}
Nicolo Cesa-Bianchi, Yoav Freund, David Haussler, David~P Helmbold, Robert~E
  Schapire, and Manfred~K Warmuth.
\newblock How to use expert advice.
\newblock \emph{Journal of the ACM (JACM)}, 1997.

\bibitem[Dud{\'i}k et~al.(2011)Dud{\'i}k, Hsu, Kale, Karampatziakis, Langford,
  Reyzin, and Zhang]{dudik2011efficient}
Miroslav Dud{\'i}k, Daniel Hsu, Satyen Kale, Nikos Karampatziakis, John
  Langford, Lev Reyzin, and Tong Zhang.
\newblock Efficient optimal learning for contextual bandits.
\newblock In \emph{Uncertainty and Artificial Intelligence (UAI)}, 2011.

\bibitem[Freund and Schapire(1997)]{Freund1997}
Yoav Freund and Robert~E Schapire.
\newblock A decision-theoretic generalization of on-line learning and an
  application to boosting.
\newblock \emph{Journal of Computer and System Sciences}, 1997.

\bibitem[Langford and Zhang(2008)]{langford2008epoch}
John Langford and Tong Zhang.
\newblock The epoch-greedy algorithm for multi-armed bandits with side
  information.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  2008.

\bibitem[Rakhlin and Sridharan(2016)]{Rakhlin2016}
Alexander Rakhlin and Karthik Sridharan.
\newblock {BISTRO:} an efficient relaxation-based method for contextual
  bandits.
\newblock In \emph{Proceedings of the Twentieth International Conference on
  Machine Learning (ICML 2016)}, 2016.
\newblock URL \url{http://arxiv.org/abs/1602.02196}.

\bibitem[Syrgkanis et~al.(2016)Syrgkanis, Krishnamurthy, and
  Schapire]{Syrgkanis2016}
Vasilis Syrgkanis, Akshay Krishnamurthy, and Robert~E. Schapire.
\newblock Efficient algorithms for adversarial contextual learning.
\newblock In \emph{Proceedings of the Twentieth International Conference on
  Machine Learning (ICML 2016)}, 2016.
\newblock URL \url{http://arxiv.org/abs/1602.02454}.

\end{thebibliography}

\documentclass{article}

\PassOptionsToPackage{numbers, compress}{natbib}
\usepackage{natbib}
\usepackage[margin=1in]{geometry}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      

\usepackage{amsmath,amsfonts,graphpap,amscd,mathrsfs,graphicx,lscape}
\newcommand*{\arginf}{\mathop{\mathrm{arginf}}}
\newcommand*{\argsup}{\mathop{\mathrm{argsup}}}
\newcommand*{\argmax}{\mathop{\mathrm{argmax}}}
\newcommand*{\argmin}{\mathop{\mathrm{argmin}}}

\usepackage{algorithmic}
\usepackage{algorithm}

\newtheorem{theorem}{Theorem}
\newtheorem{defn}{Definition}
\newtheorem{lemma}[theorem]{Lemma}

\newcommand{\Exp}[2]{{{\rm E}_{#1}\left[{#2}\right]}}

\newcommand{\dotp}[2]{\langle #1, #2 \rangle}

\newcommand{\Pols}{\Pi}
\newcommand{\pdist}{{q}}
\newcommand{\pdistt}{{\pdist_t}}

\newcommand{\cost}{{c}}
\newcommand{\ccost}{{c}}
\newcommand{\ccostt}{{\ccost_t}}

\newcommand{\qed}{\hfill\rule{2.5mm}{2.5mm}}

\newcommand{\rad}{\epsilon}
\newcommand{\vrad}{{\bf \rad}}

\newcommand{\Rel}{\textsc{Rel}}
\newcommand{\Reg}{\textsc{Reg}}

\renewcommand{\vec}[1]{\ensuremath{{\bf #1}}}

\newenvironment{rtheorem}[3][]{
\bigskip
\noindent \ifthenelse{\equal{#1}{}}{\bf #2 #3}{\bf #2 #3 (#1)}
\begin{it}
}{\end{it}}

\newcommand{\mytitle}{
Improved Regret Bounds for Oracle-Based Adversarial Contextual Bandits
}

\title{\mytitle}

\author{
Vasilis Syrgkanis\\
Microsoft Research
\and
Haipeng Luo\\
Princeton University
\and
Akshay Krishnamurthy\\
Microsoft Research
\and
Robert E. Schapire\\
Microsoft Research
}

\begin{document}

\maketitle

\begin{abstract}
  We give an oracle-based algorithm for the adversarial contextual bandit problem, where either contexts are drawn i.i.d. or the sequence of contexts is known a priori, but where the losses are picked adversarially. Our algorithm is computationally efficient, assuming access to an offline optimization oracle, and enjoys a regret of order $O((KT)^{\frac{2}{3}}(\log N)^{\frac{1}{3}})$, where $K$ is the number of actions, $T$ is the number of iterations and $N$ is the number of baseline policies. Our result is the first to break the $O(T^{\frac{3}{4}})$ barrier that is achieved by recently introduced algorithms.  Breaking this barrier was left as a major open problem. Our analysis is based on the recent relaxation based approach of~\citet{Rakhlin2016}.
\end{abstract}

\section{Introduction}

We study online decision making problems where a learner chooses an action based on some side information (context) and incurs some cost for that action with a goal of incurring minimal cost over a sequence of rounds.
These \emph{contextual online learning} settings form a powerful framework for modeling many important decision-making scenarios with applications ranging from personalized health care in the medical domain to content recommendation and targeted advertising in internet applications.
Many of these applications also involve a partial feedback component wherein costs for alternative actions are unobserved, and are typically modeled as \emph{contextual bandits}.

The contextual information present in these problems enables learning of a much richer \emph{policy} for choosing actions based on context.
In the literature, the typical goal for the learner is to have cumulative cost that is not much higher then the best policy $\pi$ in a large set $\Pi$.
This is formalized by the notion of \emph{regret}, which is the learner's cumulative cost minus the cumulative cost of the best fixed policy $\pi$ in hindsight.
Achieving this goal requires learning a rich policy, depending on $\Pi$.

Naively one can view the contextual problem as a standard online learning problem where the set of possible ``actions'' available at each iteration is the set of policies.
This perspective is fruitful, as classical algorithms, such as Hedge \cite{Freund1997,cesa1997use} and Exp4~\cite{Auer1995}, give information theoretically optimal regret bounds of $O(\sqrt{T\log(|\Pi|)})$ in full-information and $O(\sqrt{TK\log(|\Pi|)}$ in the bandit setting, where $T$ is the number of rounds, $K$ is the number of actions, and $\Pi$ is the policy set.
However, naively lifting standard online learning algorithms to the contextual setting leads to a running time that is linear in the number of policies.
Given that the optimal regret is only logarithmic in $|\Pi|$ and that our high-level goal is to learn a very rich policy, we want to capture policy classes that are exponentially large.
When we use a large policy class, existing algorithms are no longer computationally tractable.

To study this computational question, a number of recent papers have developed \emph{oracle-based algorithms} that only access the policy class through an optimization oracle for the offline full-information problem. 
Oracle-based approaches harness the research in supervised learning that focuses on designing efficient algorithms for full-information problems and uses it for online and partial-feedback problems.
Optimization oracles have been used in designing contextual bandit algorithms~\cite{Agarwal2014,langford2008epoch,dudik2011efficient} that achieve the optimal $O(\sqrt{KT\log(|\Pi|)})$ regret while also being computationally efficient (i.e. requiring $\textrm{poly}(K, \log(\Pi), T)$ oracle calls and computation).
However, these results only apply when the contexts and costs are drawn at random and identically and independently at each iteration, contrasting with the computationally inefficient approaches that can handle adversarial inputs.

Two very recent works provide the first oracle efficient algorithms for the contextual bandit problem in adversarial settings~\cite{Rakhlin2016,Syrgkanis2016}.
Rakhlin and Sridharan~\cite{Rakhlin2016} considers a setting where the contexts are drawn i.i.d. from a known distribution with adversarial costs and they provide an oracle efficient algorithm with $O(T^{\frac{3}{4}}K^{\frac{1}{2}}(\log(|\Pi|))^{\frac{1}{4}})$ regret. 
Their algorithm also applies in the transductive setting where the sequence of contexts is known a priori.
Srygkanis et. al~\cite{Syrgkanis2016} also obtain a $T^{\frac{3}{4}}$-style bound with a different oracle-efficient algorithm, but in a setting where the learner knows only the set of contexts that will arrive.
Both of these results achieve very suboptimal regret bounds, as the dependence on the number of iterations is far from the optimal $O(\sqrt{T})$-bound.
A major open question posed by both works is whether the $O(T^{\frac{3}{4}})$ barrier can be broken. 

In this paper, we provide an oracle-based contextual bandit algorithm that achieves regret $O((KT)^{\frac{2}{3}}(\log(|\Pi|))^{\frac{1}{3}})$ in both the i.i.d. context and the transductive settings considered by Rakhlin and Sridharan~\cite{Rakhlin2016}.
This bound matches that of the epoch-greedy algorithm of Langford and Zhang~\cite{langford2008epoch} that only applies to the fully stochastic setting.
As in Rakhlin and Sridharan~\cite{Rakhlin2016}, our algorithm only requires access to a \emph{value oracle}, which is weaker than the standard argmax oracle, and it makes $K+1$ oracle calls per iteration.
To our knowledge, this is the best regret bound achievable by an oracle-efficient algorithm for any adversarial contextual bandit problem.


Our algorithm and regret bound are based on a novel and intricate analysis of the minimax problem that arises in the relaxation-based framework of Rakhlin and Sridharan~\cite{Rakhlin2016}.
Our proof requires analyzing the value of a sequential game where the learner chooses a distribution over actions and then the adversary chooses a distribution over costs in some bounded finite domain, with - importantly - a bounded variance.
This is unlike the simpler minimax problem analyzed in \cite{Rakhlin2016}, where the adversary is only constrained by the range of the costs. 
We provide this tighter minimax analysis in Section~\ref{sec:regret-bound}.

Apart from showing that this more structured minimax problem has a small value, we also need to derive an oracle-based strategy for the learner that achieves the improved regret bound.
The additional constraints on the game require a much more intricate argument to derive this strategy which is an algorithm for solving a structured two-player minimax game. 
We present this part in Section \ref{sec:efficiency}.

\section{Model and Preliminaries}
\label{sec:set-up}

\paragraph{Basic notation.} Throughout the paper we denote with $x_{1:t}$ a sequence of quantities $\{x_1,\ldots,x_t\}$ and with $(x,y,z)_{1:t}$ a sequence of tuples $\{(x_1,y_1,z_1),\ldots\}$. $\emptyset$ denotes an empty sequence. The vector of ones is denoted by $\vec{1}$ and the vector of zeroes is denoted by $\vec{0}$. Denote with $[K]$ the set $\{1,\ldots,K\}$ and $\Delta_U$ the set of distributions over a set $U$. 
We also use $\Delta_K$ as a shorthand for $\Delta_{[K]}$.

\paragraph{Contextual online learning.} We consider the following version of the contextual online learning problem. On each round $t=1,\ldots,T$, the learner observes a context $x_t$ and then chooses a probability
distribution $\pdistt$ over a set of $K$ actions. The adversary then chooses a cost vector $c_t\in [0,1]^K$. The learner picks an action $\hat{y}_t$ drawn from distribution $q_t$, suffers a loss $c_t(\hat{y}_t)$ and observes only $c_t(\hat{y}_t)$ and not the loss of the other actions. 

Throughout the paper we will assume that the context $x_t$ at each iteration $t$ is drawn i.i.d. from a distribution $D$. This is referred to as the \emph{hybrid i.i.d.-adversarial} setting~\cite{Rakhlin2016}. As in prior work~\cite{Rakhlin2016}, we assume that the learner can sample contexts from this distribution as needed. It is easy to adapt the arguments in the paper to apply for the transductive setting where the learner knows the sequence of contexts that will arrive. The cost vectors $c_t$ are chosen by a non-adaptive adversary. 

The goal of the learner is to compete with a set of policies $\Pi$, where each policy $\pi\in \Pi$ is a function mapping from the set of contexts to the set of actions. The cumulative expected regret with respect to the best fixed policy in hindsight is
\[
 \Reg = \sum_{t=1}^T \pdistt\cdot \ccostt - \inf_{\pi\in\Pols}\sum_{t=1}^T \cost_t(\pi(x_t))~.
\]

\paragraph{Optimization value oracle.} We will assume that we are given access to an optimization oracle that when given as input a sequence of contexts and loss vectors $(x,c)_{1:t}$, it outputs the value of the cumulative loss of the best fixed policy: i.e.
\begin{equation}\label{eqn:oracle}
\inf_{\pi \in \Pi} \sum_{\tau=1}^t c_t(\pi(x_t))~.
\end{equation}
This can be viewed as an offline batch optimization or ERM oracle.

\subsection{Relaxation based algorithms}
We briefly review the relaxation based framework proposed in \citep{Rakhlin2016}. The reader is directed to \cite{Rakhlin2016} for a more extensive exposition. We will also slightly augment the framework by adding some internal random state that the algorithm might keep and use subsequently and which does not affect the cost of the algorithm. 

A crucial concept in the relaxation based framework is the information obtained by the learner at the end of each round $t\in [T]$, which is the following tuple:
\begin{equation*}
I_t(x_t,q_t,\hat{y}_t,c_t,S_t)=(x_t,q_t,\hat{y}_t,c_t(\hat{y}_t),S_t)
\end{equation*}
where $\hat{y}_t$ is the realized chosen action drawn from the distribution $q_t$ and $S_t$ is some random string drawn from some distribution that can depend on $q_t$, $\hat{y}_t$ and $c_t(\hat{y}_t)$ and which can be used by the algorithm in subsequent rounds.

\begin{defn} A partial-information relaxation $\Rel(\cdot)$ is a function that maps $(I_1,\ldots,I_t)$ to a real value for any $t\in [T]$. A partial-information relaxation is admissible if for any $t\in [T]$, and for all $I_1,\ldots,I_{t-1}$:
\begin{equation}\label{eqn:cond1}
\Exp{x_t}{\inf_{q_t}\sup_{c_t}\Exp{\hat{y}_t\sim q_t, S_t}{c_t(\hat{y}_t)+\Rel(I_{1:t-1},I_t(x_t,q_t,\hat{y}_t,c_t,S_t))}}\leq \Rel(I_{1:t-1})
\end{equation}
and for all $x_{1:T},c_{1:T}$ and $q_{1:T}$:
\begin{equation}\label{eqn:cond2}
\Exp{\hat{y}_{1:T}\sim q_{1:T},S_{1:T}}{\Rel(I_{1:T})}\geq -\inf_{\pi\in \Pi}\sum_{t=1}^T c_t(\pi(x_t))~.
\end{equation}
\end{defn}

\begin{defn} Any randomized strategy $q_{1:T}$ that certifies inequalities \eqref{eqn:cond1} and \eqref{eqn:cond2} is called an admissible strategy.
\end{defn}

A basic lemma proven in \cite{Rakhlin2016} is that if one constructs a relaxation and a corresponding admissible strategy, then the expected regret of the admissible strategy is upper bounded by the value of the relaxation at the beginning of time. 

\begin{lemma}[\cite{Rakhlin2016}]\label{lem:regret-bound} Let $\Rel$ be an admissible relaxation and $q_{1:T}$ be an admissible strategy. Then for any $c_{1:T}$,
we have
\begin{equation*}
\Exp{}{\Reg}\leq \Rel(\emptyset)~.
\end{equation*}
\end{lemma}

We will utilize this framework and construct a novel relaxation with an admissible strategy. We will show that the value of the relaxation at the beginning of time is upper bounded by the desired improved regret bound and that the admissible strategy can be efficiently computed assuming access to an optimization value oracle.

\section{A Faster Contextual Bandit Algorithm}\label{sec:regret-bound}
First we define an unbiased estimator for each loss vector $c_t$.
In addition to doing the usual importance weighting, we also discretize the estimated loss to either $0$ or $L$ for some constant $L \geq K$ to be specified later.
Specifically, consider a random variable $X_t$ which we construct at the end of each iteration conditioning on $\hat{y}_t$:
\begin{align}
	X_t = \begin{cases}
	1 & \text{ with probability } \frac{c_t(\hat{y}_t)}{L q_t(\hat{y}_t)} ,\\
	0 & \text{ with the remainig probability } .
	\end{cases}\label{eqn:rv_X}
\end{align}
This is a valid random variable whenever $\min_i q_t(i)\geq \frac{1}{L}$, which will be ensured by the algorithm. This is the only random variable in the random string $S_t$ that we used in the general formulation of the relaxation framework.

Now the construction of an unbiased estimate for each $c_t$ based on the information $I_t$ collected at the end of each round is: $\hat{c}_t = L  X_t \vec{e}_{\hat{y}_t}$. Observe that for any $i\in [K]$:
\begin{equation*}
\Exp{\hat{y}_t\sim q_t, X_t}{\hat{c}_t(i)} = L\cdot \Pr[\hat{y}_t=i]\cdot\Pr[X_t=1|\hat{y}_t=i] = L \cdot q_t(i) \cdot \frac{c_t(i)}{L q_t(i)} = c_t(i)~.
\end{equation*}
Hence, $\hat{c}_t$ is an unbiased estimate of $c_t$.

We are now ready to define our relaxation. Let $\vrad_t\in \{-1,1\}^K$ be a Rademacher random vector (i.e. each coordinate is an independent Rademacher random variable, which is $-1$ or $1$ with equal probability), and let $Z_t\in \{0,L\}$ be a random variable which is $L$ with probability $K/L$ and $0$ otherwise. 
With the notation $\rho_t = (x,\epsilon,Z)_{t+1:T}$, our relaxation is defined as follows:
\begin{equation}
\Rel(I_{1:t}) = \Exp{\rho_t}{R((x,\hat{c})_{1:t},\rho_t)}\label{defn:relaxation}~, 
\end{equation}
where
\[ R((x,\hat{c})_{1:t},\rho_t) = -\inf_{\pi \in \Pols} \left(\sum_{\tau=1}^t \hat{\cost}_{\tau}(\pi(x_\tau)) + \sum_{\tau=t+1}^T 2\rad_{\tau}(\pi(x_\tau))Z_t\right) + (T-t)K/L~. \]
Note that $\Rel(\emptyset)$ is the following quantity, whose first part resembles a Rademacher average:
\begin{equation*}
{\cal R}_{\Pols} = 2\Exp{(x,\vrad,Z)_{1:T}}{\inf_{\pi \in \Pols} \sum_{\tau=t+1}^T \rad_{\tau}(\pi(x_\tau))Z_t}+TK/L~.
\end{equation*}
Using the following Lemma (whose proof is deferred to the supplementary material) and the fact $\Exp{}{Z_t^2} \leq KL$, we can upper bound ${\cal R}_{\Pols}$ by $O(\sqrt{TKL\log(N)}+TK/L)$,
which after tuning $L$ will give the claimed $O(T^{2/3})$ bound.

\begin{lemma}\label{lem:rademacher-bound}
Let $\epsilon_{t}$ be Rademacher random vectors, and $Z_t$ be non-negative real-valued random variables, such that $\Exp{}{Z_t^2}\leq M$. Then:
\begin{equation*}
\Exp{Z_{1:T},\epsilon_{1:T}}{\sup_{\pi\in \Pi} \sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}\leq \sqrt{2T M \log(N)}
\end{equation*}
\end{lemma}

To show an admissible strategy for our relaxation, we next introduce some more notation.  Let 
\begin{equation*}
D = \{L\cdot\vec{e}_i: i\in [K]\}\cup \{\vec{0}\},
\end{equation*}
where $(\vec{e}_1,\ldots,\vec{e}_K)$ are the orthonormal basis vectors (i.e. $\vec{e}_i$ is $1$ at coordinate $i$ and zero otherwise), and $\vec{0}$ is the all zeros vector. We will denote with $\Delta_D$ the set of distributions over $D$. For a distribution $p\in \Delta(D)$, we will denote with $p(i)$, for $i\in \{0,\ldots,K\}$, the probability assigned to vector $\vec{e}_i$, with the convention that $\vec{e}_0=\vec{0}$.
Also let $\Delta'_D = \{ p \in \Delta_D : p(i) \leq 1/L,  \forall i \in [K] \}$.

Based on this notation our admissible strategy is defined as 
\begin{equation}\label{eqn:strategy}
q_t = \Exp{\rho_t}{q_t(\rho_t)}    \text{\quad where \quad} q_t(\rho_t) = \left(1-\frac{K}{L}\right) q_t^*(\rho_t)+\frac{1}{L}{\bf 1} 
\end{equation}
and
\begin{equation}\label{eqn:distt}
q_t^*(\rho_t) = \argmin_{q\in \Delta_K}\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t}{\dotp{q}{\hat{c}_t} + R((x,\hat{c})_{1:t}, \rho_t)}~.
\end{equation}
Algorithm \ref{alg:cont-experts} implements this admissible strategy.
Note that it suffices to use $q_t(\rho_t)$ instead of $q_t$ for a random draw $\rho_t$ in the algorithm to ensure the exact same guarantee in expectation.
Moreover, in Section~\ref{sec:efficiency} we will show that $q_t(\rho_t)$ can be computed efficiently using an optimization value oracle.

We now prove that our relaxation and strategy are indeed admissible.

\begin{algorithm}[t]
  \caption{A New Contextual Bandit Algorithm}
\label{alg:cont-experts}
\begin{algorithmic}
	\STATE {\bfseries Input:} parameter $L \geq K$
	\FOR{each time step $t\in [T]$}
   	\STATE Observe $x_t$. Draw $\rho_t = (x,\epsilon,Z)_{t+1:T}$ where each $x_\tau$ is drawn from the distribution of contexts, $\epsilon_\tau$ is a Rademacher random vectors and $Z_\tau \in \{0,L\}$ is $L$ with probability $K/L$ and $0$ otherwise.
	\STATE Compute $q_t(\rho_t)$ based on Eq.~\eqref{eqn:strategy} (using Algorithm~\ref{alg:q_t}).
	\STATE Predict $\hat{y}_t\sim q_t(\rho_t)$ and observe $c_t(\hat{y}_t)$.
	\STATE Create an estimate $\hat{c}_t = L X_t \vec{e}_{\hat{y}_t}$, where $X_t$ is defined in Eq.~\eqref{eqn:rv_X} with $q_t$ in that equation instantiated with $q_t(\rho_t)$.
	\ENDFOR
\end{algorithmic}
\end{algorithm}

\begin{theorem} The relaxation defined in Equation \eqref{defn:relaxation} is admissible. An admissible randomized strategy for this relaxation is given by \eqref{eqn:strategy}. The expected regret of the Algorithm~\ref{alg:cont-experts} is upper bounded by 
\begin{equation}\label{eqn:regret}
2\sqrt{2TKL\log(N)}+TK/L,
\end{equation}
for any $L \geq K$. 
Specifically, setting $L = \left( KT/\log(N) \right)^{\frac{1}{3}}$ when $T \geq K^2\log(N)$, the regret is of order $O((KT)^{\frac{2}{3}}(\log(N))^{\frac{1}{3}})$.
\end{theorem}
\begin{proof}
We verify the two conditions for admissibility.
\paragraph{Final condition.}
It is clear that inequality \eqref{eqn:cond2} is satisfied since $\hat{c}_t$ are unbiased estimates of $c_t$:
\begin{align*}
\Exp{\hat{y}_{1:T}, X_{1:T}}{\Rel(I_{1:T})} =~& \Exp{\hat{y}_{1:T}, X_{1:T}}{\sup_{\pi \in \Pols}- \sum_{\tau=1}^T \hat{\cost}_{\tau}(\pi(x_\tau))} \\
\geq~& \sup_{\pi \in \Pols}-\Exp{\hat{y}_{1:T}, X_{1:T}}{ \sum_{\tau=1}^T \hat{\cost}_{\tau}(\pi(x_\tau))} 
=  \sup_{\pi \in \Pols}- \sum_{\tau=1}^T \cost_{\tau}(\pi(x_\tau))
\end{align*}

\paragraph{$t$-th Step condition.} We now check that inequality \eqref{eqn:cond1} is also satisfied at some time step $t\in [T]$. We reason conditionally on the observed context $x_t$ and show that $q_t$ defines an admissible strategy for the relaxation.
Let $q_t^* = \Exp{\rho_t}{q_t^*(\rho_t)} $. First observe that:
\begin{equation*}
\Exp{\hat{y}_t, X_t}{c_t(\hat{y}_t)} =\Exp{\hat{y}_t\sim q_t}{c_t(\hat{y}_t)}=  \dotp{q_t}{c_t}\leq \dotp{q_t^*}{c_t}+\frac{1}{L}\dotp{{\bf 1}}{c_t}\leq \Exp{\hat{y}_t, X_t}{\dotp{q_t^*}{\hat{c}_t}}+\frac{K}{L}
\end{equation*}
We remind that $\hat{c}_t=L X_t \vec{e}_{\hat{y}_t}$ is the unbiased estimate, which is a deterministic function of $\hat{y}_t$ and $X_t$.

Hence:
\begin{align*}
\sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t, X_t}{c_t(\hat{y}_t)+\Rel(I_{1:t})}\leq  \sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t, X_t}{\dotp{q_t^*}{\hat{c}_t}+\Rel(I_{1:t})}+\frac{K}{L}
\end{align*}
We now work with the first term of the right hand side:
\begin{align*}
\sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t, X_t}{\dotp{q_t^*}{\hat{c}_t}+\Rel(I_{1:t})}=~& \sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t, X_t}{\dotp{q_t^*}{\hat{c}_t}+\Exp{\rho_t}{R((x,\hat{c})_{1:t},\rho_t}}\\
=~&\sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t, X_t}{\Exp{\rho_t}{\dotp{q_t^*(\rho_t)}{\hat{c}_t}+R((x,\hat{c})_{1:t},\rho_t)}}
\end{align*}
Observe that $\hat{c}_t$ is a random variable taking values in $D$ and such that the probability that it is equal to $L\vec{e}_i$ can be upper bounded as:
\begin{equation*}
\Pr[\hat{c}_t = L\vec{e}_i] = \Exp{\rho_t}{\Pr[\hat{c}_t = L\vec{e}_i|\rho_t]}= \Exp{\rho_t}{q_t(\rho_t)(i) \frac{c_t(i)}{L\cdot q_t(\rho_t)(i)}}\leq 1/L.
\end{equation*}
Thus we can upper bound the latter quantity by the supremum over all distributions in $\Delta'_D$, i.e.:
\begin{align*}
\sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t,X_t}{\dotp{q_t^*}{\hat{c}_t}+\Rel(I_{1:t})}\leq~& \sup_{p_t\in \Delta'_D} \Exp{\hat{c}_t\sim p_t}{\Exp{\rho_t}{\dotp{q_t^*(\rho)}{\hat{c}_t}+R((x,\hat{c})_{1:t},\rho_t)}}
\end{align*}
Now we can continue by pushing the expectation over $\rho_t$ outside of the supremum, i.e.
\begin{align*}
\sup_{c_t\in [0,1]^K} \Exp{\hat{y}_t\sim q_t,X_t}{\dotp{q_t^*}{\hat{c}_t}+\Rel(I_{1:t})}\leq~& \Exp{\rho_t}{\sup_{p_t\in \Delta'_D} \Exp{\hat{c}_t\sim p_t}{\dotp{q_t^*(\rho_t)}{\hat{c}_t}+R((x,\hat{c})_{1:t},\rho_t)}}
\end{align*}
and working conditionally on $\rho_t$. Observe that by the definition of $q_t^*(\rho_t)$ the quantity inside the expectation is equal to:
\begin{equation*}
\inf_{q\in \Delta_K}\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t}{\dotp{q}{\hat{c}_t} + R((x,\hat{c})_{1:t},\rho_t)}
\end{equation*}
We can now apply the minimax theorem and upper bound the above by:
\begin{equation*}
\sup_{p_t \in \Delta'_D}\inf_{q\in \Delta_K}\Exp{\hat{c}_t\sim p_t}{\dotp{q}{\hat{c}_t} + R((x,\hat{c})_{1:t},\rho_t)}
\end{equation*}
Since the inner objective is linear in $q$, we continue with
\begin{equation*}
\sup_{p_t \in \Delta'_D}\min_{i}\Exp{\hat{c}_t\sim p_t}{\hat{c}_t(i) + R((x,\hat{c})_{1:t},\rho_t)}
\end{equation*}
We can now expand the definition of $R(\cdot)$:
\begin{equation*}
\sup_{p_t \in \Delta'_D}\min_{i}\Exp{\hat{c}_t\sim p_t}{\hat{c}_t(i) + \sup_{\pi \in \Pi}-\left(\sum_{\tau=1}^t \hat{\cost}_{\tau}(\pi(x_\tau)) + \sum_{\tau=t+1}^T 2\rad_{\tau}(\pi(x_\tau))Z_t\right)} + (T-t)K/L
\end{equation*}
With the notation
\begin{equation*}
A_{\pi} = -\sum_{\tau=1}^{t-1} \hat{\cost}_{\tau}(\pi(x_\tau)) - \sum_{\tau=t+1}^T 2\rad_{\tau}(\pi(x_\tau))Z_t
\end{equation*}
we re-write the above quantity as:
\begin{equation*}
\sup_{p_t \in \Delta'_D}\min_{i}\Exp{\hat{c}_t\sim p_t}{\hat{c}_t(i) + \sup_{\pi \in \Pi}(A_{\pi}-\hat{c}_t(\pi(x_t)))} + (T-t)K/L
\end{equation*}
We now upper bound the first term. The extra term $(T-t)K/L$ will be combined with the extra $K/L$ that we have abandoned to give the correct term $(T-(t-1))K/L$ needed for $\Rel(I_{1:t-1})$. 

Observe that we can re-write the first term by using symmetrization as:
\begin{align*}
&\sup_{p_t \in \Delta'_D}\min_{i}\Exp{\hat{c}_t\sim p_t}{\hat{c}_t(i) + \sup_{\pi \in \Pi}(A_{\pi}-\hat{c}_t(\pi(x_t)))}\\
&= \sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t}{\sup_{\pi \in \Pi}(A_{\pi}+\min_{i} \Exp{\hat{c}_t'\sim p_t}{\hat{c}_t'(i)}-\hat{c}_t(\pi(x_t)))}\\
&\leq \sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t}{\sup_{\pi \in \Pi}(A_{\pi}+\Exp{\hat{c}_t'\sim p_t}{\hat{c}_t'(\pi(x_t))}-\hat{c}_t(\pi(x_t)))}\\
&\leq \sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t,\hat{c}_t'\sim p_t}{\sup_{\pi \in \Pi}(A_{\pi}+\hat{c}_t'(\pi(x_t))-\hat{c}_t(\pi(x_t)))}\\
&=\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t,\hat{c}_t'\sim p_t, \delta}{\sup_{\pi \in \Pi}(A_{\pi}+\delta\left(\hat{c}_t'(\pi(x_t))-\hat{c}_t(\pi(x_t)))\right)}\\
&\leq\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t, \delta}{\sup_{\pi \in \Pi}(A_{\pi}+2\delta\hat{c}_t(\pi(x_t)))}\\
\end{align*}
where $\delta$ is a random variable which is $-1$ and $1$ with equal probability. The last inequality follows by splitting the supremum into two equal parts.

Conditioning on $\hat{c}_t$, consider the random variable $M_t$ which is $-\max_i \hat{c}_t(i)$ or $\max_i \hat{c}_t(i)$ on the coordinates where $\hat{c}_t$ is equal to zero and equal to $\hat{c}_t$ on the coordinate that achieves the maximum. This is clearly an unbiased estimate of $\hat{c}_t$. Thus we can upper bound the last quantity by:
\begin{align*}
\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t, \delta}{\sup_{\pi \in \Pi}(A_{\pi}+2\delta\Exp{}{M_t(\pi(x_t))|\hat{c}_t})}
\leq
\sup_{p_t \in \Delta'_D}\Exp{\hat{c}_t\sim p_t, \delta,M_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\delta M_t(\pi(x_t)))}
\end{align*}
The random vector $\delta M_t$, conditioning on $\hat{c}_t$, is equal to $-\max_i \hat{c}_t(i)$ or $\max_i \hat{c}_t(i)$ with equal probability independently on each coordinate. Moreover, observe that for any distribution $p_t\in \Delta'_D$, the distribution of the maximum coordinate of $\hat{c}_t$ has support on $\{0,L\}$ and is equal to $L$ with probability at most $K/L$. Since the objective only depends on the distribution of the maximum coordinate of $\hat{c}_t$, we can continue the upper bound with a supremum over any distribution of random vectors whose coordinates are $0$ with probability at least $1-K/L$ and otherwise are $-L$ or $L$ with equal probability. Specifically, let $\epsilon_t$ be a Rademacher random vector, we continue with:
\begin{align*}
\sup_{Z_t\in \Delta_{\{0,L\}}:Pr[Z_t=L]\leq K/L}\Exp{\epsilon_t,Z_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\epsilon_t(\pi(x_t))Z_t)}
\end{align*}
Now observe that if we denote with $a=\Pr[Z_t=L]$, the above is equal to:
\begin{align*}
\sup_{a:0\leq a \leq K/L}\left((1-a)\sup_{\pi \in \Pi}(A_{\pi})+a\Exp{\epsilon_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\epsilon_t(\pi(x_t))L)}\right)
\end{align*}
We now argue that this supremum is achieved by setting $a=K/L$. For that it suffices to show that:
\begin{equation*}
\sup_{\pi \in \Pi}(A_{\pi}) \leq\Exp{\epsilon_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\epsilon_t(\pi(x_t))L)}
\end{equation*}
which is true by observing that with $\pi^*=\argsup_{\pi \in \Pi}(A_{\pi})$ one has:
\begin{equation*}
\Exp{\epsilon_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\epsilon_t(\pi(x_t))L)}\geq \Exp{\epsilon_t}{A_{\pi^*}+2\epsilon_t(\pi^*(x_t))L)}=
A_{\pi^*}+\Exp{\epsilon_t}{2\epsilon_t(\pi^*(x_t))L)} = A_{\pi^*}~.
\end{equation*}

Thus we can upper bound the quantity we want by:
\begin{equation*}
\Exp{\epsilon_t,Z_t}{\sup_{\pi \in \Pi}(A_{\pi}+2\epsilon_t(\pi(x_t))Z_t}
\end{equation*}
where $\epsilon_t$ is a Rademacher random vector and $Z_t$ is now a random variable which is equal to $L$ with probability $K/L$ and is equal to $0$ with the remaining probability. 

Taking expectation over $\rho_t$ and $x_t$ and adding the $(T-(t-1))K/L$ term that we abandoned, we arrive at the desired upper bound of $\Rel(I_{1:t-1})$. This concludes the proof of admissibility.

\paragraph{Regret bound.} By applying Lemma \ref{lem:rademacher-bound} with $E[Z_t^2] = L^2 \Pr[Z_t=L]= KL$ and invoking Lemma \ref{lem:regret-bound}, we get the regret bound in Equation \eqref{eqn:regret}.
\qed\end{proof}

\section{Computational Efficiency}\label{sec:efficiency}

In this section we will argue that if one is given access to a value optimization oracle~\eqref{eqn:oracle}, then one can run Algorithm \ref{alg:cont-experts} efficiently. Specifically, we will show that the minimizer of Equation \eqref{eqn:distt} can be computed efficiently (see Algorithm~\ref{alg:q_t}).

\begin{algorithm}[t]
  \caption{Computing $q^*_t(\rho_t)$}
\label{alg:q_t}
\begin{algorithmic}
	\STATE {\bfseries Input:} a value optimization oracle, $(x, \hat{c})_{1:t-1}$, $x_t$ and $\rho_t$.
	\STATE {\bfseries Output:} $q \in \Delta_K$ as a solution of Eq.~\eqref{eqn:distt}.
	\STATE ~
	\STATE Compute $\psi_i$ as in Eq.~\eqref{eqn:psi} for all $i = 0, \ldots, K$ using the optimization oracle.
	\STATE Compute $\phi_i = \frac{\psi_i-\psi_0}{L}$ for all $i \in [K]$.
	\STATE Let $m = 1$ and $q = \vec{0}$.
	\FOR{each coordinate $i \in [K]$}
		\STATE Set $q(i) = \min\{(\phi_i)^+, m\}$.
		\STATE Update $m \leftarrow m - q(i)$.			
	\ENDFOR
	\STATE Distribute $m$ arbitrarily on the coordinates of $q$ if $m > 0$.	
\end{algorithmic}
\end{algorithm}

\begin{lemma}
Computing the quantity defined in equation \eqref{eqn:distt} for any given $\rho_t$ can be done in time $O(K)$ and with only $K+1$ accesses to a value optimization oracle. 
\end{lemma}
\begin{proof}
For $i\in \{0,\ldots,K\}$, let:
\begin{equation}\label{eqn:psi}
\psi_i = \inf_{\pi\in \Pi} \left(\sum_{\tau=1}^{t-1} \hat{\cost}_{\tau}(\pi(x_\tau))+ L \vec{e}_i(\pi(x_t)) + \sum_{\tau=t+1}^T 2\rad_{\tau}(\pi(x_\tau))Z_t\right) 
\end{equation}
with the convention $\vec{e}_0 = \vec{0}$. Then observe that we can re-write the definition of $q_t^*(\rho_t)$ as:
\begin{equation*}
q_t^*(\rho_t) = \arginf_{q\in \Delta_K} \sup_{p_t \in \Delta'_D} \sum_{i=1}^K p_t(i)(L\cdot q(i) - \psi_i)- p_t(0) \cdot  \psi_0
\end{equation*}

Observe that each $\psi_i$ can be computed with a single oracle access. Thus we can assume that all $K+1$ $\psi$'s are computed efficiently and are given. We now argue how to compute the minimizer.

For each given $q$, the supremum over $p_t$ can be characterized as follows. With the notation $z_i=L\cdot q(i)-\psi_i$ and $z_0=-\psi_0$ we re-write the minimax quantity as: 
\begin{equation*}
q_t^*(\rho_t) = \arginf_{q\in \Delta_K} \sup_{p_t \in \Delta'_D} \sum_{i=1}^K p_t(i)\cdot z_i + p_t(0) \cdot z_0
\end{equation*}
Observe that if we didn't have the constraint that $p_t(i)\leq 1/L$ for $i>0$, then we would have put all the probability mass on the maximum of the $z_i$. However, now that we are constrained we will simply put as much probability mass as allowed on the maximum coordinate $\argmax_{i\in \{0,\ldots,K\}} z_i$ and continue to the next highest quantity. We repeat this until reaching the quantity $z_0$. At that point, the probability mass that we can put on coordinate $0$ is unconstrained. Thus we can put all the remaining probability mass on this coordinate. 

Let $z_{(1)},z_{(2)},\ldots,z_{(K)}$ denote the ordered $z_i$ quantities for $i>0$ (from largest to smallest). Moreover, let $\mu\in [K]$ be the largest index such that $z_{(\mu)}\geq z_0$. By the above reasoning we get that for a given $q$, the supremum over $p_t$ is equal to:
\begin{equation*}
\sum_{t=1}^{\mu} \frac{z_{(t)}}{L}+\left(1-\frac{\mu}{L}\right)z_0=
\sum_{t=1}^{\mu} \frac{z_{(t)}-z_0}{L} + z_0~.
\end{equation*}
Now since for any $t>\mu$, $z_{(t)}<z_0$, we can write the latter as:
\begin{equation*}
\sum_{t=1}^{\mu} \frac{z_{(t)}}{L}+\left(1-\frac{\mu}{L}\right)z_0=
\sum_{i=1}^{K} \frac{(z_{i}-z_0)^+}{L} + z_0
\end{equation*}
with the convention $(x)^+ = \max\{ x, 0\}$. We thus further re-write the minimax expression as:
\begin{equation*}
q_t^*(\rho_t) = \arginf_{q\in \Delta_K} \sum_{i=1}^{K} \frac{(z_{i}-z_0)^+}{L} + z_0 =  \arginf_{q\in \Delta_K} \sum_{i=1}^{K} \frac{(z_{i}-z_0)^+}{L}=
 \arginf_{q\in \Delta_K} \sum_{i=1}^{K} \left(q(i)-\frac{\psi_i-\psi_0}{L}\right)^+
\end{equation*}
Let $\phi_i = \frac{\psi_i-\psi_0}{L}$. The expression becomes: $q_t^*(\rho_t) =\arginf_{q\in \Delta_K}\sum_{t=1}^{K} (q(i)-\phi_i)^+$.

The latter is minimized as follows: consider any $i\in [K]$ such that $\phi_i\leq 0$. Then putting any positive mass on such a coordinate $i$ is going to lead to a marginal increase of $1$. On the other hand if we put some mass on an index $\phi_i>0$, then that will not increase the objective until we reach the point where $q(i)=\phi_i$. Thus a minimizer will distribute probability mass of $\min\{\sum_{i:\phi_i>0}\phi_i,1\}$, on the coordinates for which $\phi_i>0$. The remainder mass (if any) can be distributed arbitrarily. See Algorithm~\ref{alg:q_t} for details.
%
\qed\end{proof}

\section{Discussion}\label{sec:discussion}
In this paper, we present a new oracle-efficient algorithm for adversarial contextual bandits and we prove that it achieves $O((KT)^{2/3}\log(|\Pi|)^{1/3})$ regret in the settings studied by Rakhlin and Sridharan~\cite{Rakhlin2016}.
This is the best regret bound that we are aware of among oracle-based algorithms.

While our bound improves on the $O(T^{3/4})$ bounds in prior
work~\cite{Rakhlin2016,Syrgkanis2016}, achieving the optimal
$O(\sqrt{TK\log(|\Pi|)})$ regret bound with an oracle based approach
still remains an important open question.
Another interesting avenue for future work involves understanding the role of transductivity assumptions and developing an algorithm that can handle the non-transductive fully adversarial setting.
We look forward to pursuing these directions.

\begin{small}
\bibliographystyle{plainnat}
\bibliography{poa_survey}
\end{small}

\newpage

\appendix
\setcounter{page}{1}

\begin{center}
\bf \Large Supplementary material for \\ ``\mytitle''
\end{center}

\section{Supplementary Lemma}

\begin{rtheorem}{Lemma}{\ref{lem:rademacher-bound}.}
Let $\epsilon_{t}$ be Rademacher random vectors, and $Z_t$ be non-negative real-valued random variables, such that $\Exp{}{Z_t^2}\leq M$. Then:
\begin{equation*}
\Exp{Z_{1:T},\epsilon_{1:T}}{\sup_{\pi\in \Pi} \sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}\leq \sqrt{2T M \log(N)}
\end{equation*}
\end{rtheorem}
\begin{proof}
\begin{align*}
\Exp{Z_{1:T},\epsilon_{1:T}}{\sup_{\pi\in \Pi} \sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t} =~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\Exp{\epsilon_{1:T}}{\log\left(\sup_{\pi\in \Pi} e^{\lambda\sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}\right)}}\\
\leq~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(\Exp{\epsilon_{1:T}}{\sup_{\pi\in \Pi} e^{\lambda\sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}}\right)}\\
\leq~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(\Exp{\epsilon_{1:T}}{\sum_{\pi\in \Pi} e^{\lambda\sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}}\right)}\\
=~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(\sum_{\pi\in \Pi} \Exp{\epsilon_{1:T}}{\prod_{t=1}^T e^{\lambda \epsilon_t(\pi(x_t))\cdot Z_t}}\right)}\\
=~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(\sum_{\pi\in \Pi} \prod_{t=1}^T\Exp{\epsilon_{t}}{ e^{\lambda \epsilon_t(\pi(x_t))\cdot Z_t}}\right)}
\end{align*}
Now observe that $\Exp{\epsilon_{t}}{ e^{\lambda \epsilon_t(\pi(x_t))\cdot Z_t}} = \frac{e^{\lambda \cdot Z_t}+e^{-\lambda \cdot Z_t}}{2}\leq e^{\lambda^2\cdot Z_t^2/2}$. Thus:
\begin{align*}
\Exp{Z_{1:T},\epsilon_{1:T}}{\sup_{\pi\in \Pi} \sum_{t=1}^T \epsilon_t(\pi(x_t))\cdot Z_t}
\leq~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(\sum_{\pi\in \Pi} \prod_{t=1}^T e^{\lambda^2\cdot Z_t^2/2}\right)}\\
=~&
\Exp{Z_{1:T}}{\frac{1}{\lambda}\log\left(N e^{\lambda^2\sum_{t=1}^T Z_t^2/2}\right)}\\
=~&
\frac{1}{\lambda}\log(N)+\lambda \Exp{Z_{1:T}}{\sum_{t=1}^T Z_t^2/2}\\
\leq~&
\frac{1}{\lambda}\log(N)+\lambda MT/2
\end{align*}
Optimizing over $\lambda$ yields the result.
\qed\end{proof}

\end{document}